The molar mass of platinum is 195 g/mol. NA = 6.02 x 1023 units/mol

a- Find the number of moles present in 3.2 g of a platinum sample.

b- What is the mass of 2.75 x 10 24 atoms of platinum?


1
Expert's answer
2021-06-11T01:58:09-0400

Solution


a.) No of moles=

3.2g of platinum"\times \frac{1 mol}{195g/mol}"

=0.0164 moles

b.)

2.75 "\times 10^{24}" atoms"\times \frac{1mole}{6.02\times10^{24}}"

=4.568 moles of platinum

Mass= "4.568 moles\times\frac{195g}{1mole}"

=890.7g of platinum